# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Acknowledgements.
Appendix A Derivations of equations
We derive Eqs. (11), (12), and (14) for the general framework in Sec. II.2, properties of the weight $w_{mn}^{(K)}(\lambda)$ of the ground-state evolution in Sec. III.1, and the partial action decomposition in Sec. V.1. Throughout this section, we fix $\lambda$ and omit the $\lambda$ dependencies of quantities such as $H(\lambda)$, $\epsilon_{n}(\lambda)$, $|\phi_{n}(\lambda)\rangle$, $\mathcal{S}_{\lambda}^{(\mathcal{P})}[V]$, $\Phi(\lambda)$, $\mathcal{P}_{\lambda}(x)$, $w_{mn}^{(K)}(\lambda)$, $w_{n}^{(K)}(\lambda)$, and $E_{\lambda}^{(K)}$ for notational simplicity.
Algebraic characterizations
We derive the algebraic characterization of the exact AGP in Eq. (11). We first calculate $\partial^{\prime}_{\lambda}\mathcal{P}(H)$. The derivative of the projector $|\phi_{n}\rangle\langle\phi_{n}|$ is
$\partial_{\lambda}\bigl{[}|\phi_{n}\rangle\langle\phi_{n}|\bigr{]}$
$=|\partial_{\lambda}\phi_{n}\rangle\langle\phi_{n}|+|\phi_{n}\rangle\langle\partial_{\lambda}\phi_{n}|$
$=\sum_{m}|\phi_{m}\rangle\langle\phi_{m}|\partial_{\lambda}\phi_{n}\rangle\langle\phi_{n}|+\sum_{m}|\phi_{n}\rangle\langle\partial_{\lambda}\phi_{n}|\phi_{m}\rangle\langle\phi_{m}|$
$=-i\hbar^{-1}\sum_{m}\bigl{(}\Phi_{mn}|\phi_{m}\rangle\langle\phi_{n}|-\Phi_{nm}|\phi_{n}\rangle\langle\phi_{m}|\bigr{)},$ (36)
where the last equality uses the definition of $\Phi$ in Eq. (4) and $\langle\partial_{\lambda}\phi_{n}|\phi_{m}\rangle+\langle\phi_{n}|\partial_{\lambda}\phi_{m}\rangle=\partial_{\lambda}[\langle\phi_{n}|\phi_{m}\rangle]=0$. Using Eq. (36), we can calculate the $\lambda$-derivative of the fictitious Hamiltonian $\mathcal{P}(H)=\sum_{n}\mathcal{P}(\epsilon_{n})|\phi_{n}\rangle\langle\phi_{n}|$ as
@@ Appendix A Derivations of equations
$\quad-i\hbar^{-1}\sum_{nm}\mathcal{P}(\epsilon_{n})\bigl{(}\Phi_{mn}|\phi_{m}\rangle\langle\phi_{n}|-\Phi_{nm}|\phi_{n}\rangle\langle\phi_{m}|\bigr{)}$
@@
$\quad+i\hbar^{-1}\sum_{nm}\bigl{[}\mathcal{P}(\epsilon_{m})-\mathcal{P}(\epsilon_{n})\bigr{]}\Phi_{mn}|\phi_{m}\rangle\langle\phi_{n}|$
@@
Variational formula
We show that the algebraic characterization in Eq. (11) is equivalent to the variational expression in Eq. (12), which extends the derivation of the conventional variational method in Ref. [27, 8]. The algebraic characterization says that $\partial^{\prime}_{\lambda}\mathcal{P}(H)-i\hbar^{-1}[\mathcal{P}(H),\Phi]$ is diagonal with respect to the Hamiltonian basis. The diagonal elements come solely from $\partial^{\prime}_{\lambda}\mathcal{P}(H)$ since $[\mathcal{P}(H),\Phi]$ has no diagonal elements. Let $\Psi\coloneqq\sum_{n}|\phi_{n}\rangle\langle\phi_{n}|\partial^{\prime}_{\lambda}\mathcal{P}(H)|\phi_{n}\rangle\langle\phi_{n}|$ denote the diagonal part of $\partial^{\prime}_{\lambda}\mathcal{P}(H)$. Then, the algebraic characterization is equivalent to
@@
Expansion in eigenbasis
We derive the expansion of $\mathcal{S}^{(\mathcal{P})}[V]$ in Eq. (14). Using the expression of $\partial^{\prime}_{\lambda}\mathcal{P}(H)$ in Eq. (37), we can rewrite the operator inside the Hilbert–Schmidt norm as
@@
Since the Hilbert–Schmidt norm equals the sum of the squared absolute values of the elements, the action is given by
@@
$\quad+\hbar^{-2}\sum_{nm}\bigl{[}\mathcal{P}_{\lambda}(\epsilon_{m})-\mathcal{P}_{\lambda}(\epsilon_{n})\bigr{]}^{2}|\Phi_{mn}-V_{mn}|^{2},$ (42)
which reproduces Eq. (14).
@@
Properties of the weight for ground-state evolution
We derive the properties of $w_{mn}^{(K)}$ for ground-state evolution in Sec. III.1. First, we derive the explicit expression of $w_{mn}^{(K)}$ in Eq. (23) as
@@
$=\sum_{s=0}^{K-1}\sum_{u=0}^{K-1}(\epsilon_{n}-E^{(K)})^{2K-2-(s+u)}(\epsilon_{m}-E^{(K)})^{s+u}$
$=\!\sum_{\smash{s=-(K-1)}}^{K-1}\!(K-|s|)(\epsilon_{n}-E^{(K)})^{K-1-s}(\epsilon_{m}-E^{(K)})^{K-1+s}.$ (43)
@@
Combining with above, we obtain

### v3

@@ Acknowledgements.
Data Availability
The data that support the findings of this article are openly available [73].
Appendix A Appendix to the general framework
Locality of the conventional action
We show that the conventional action is given by a sum of local terms, assuming that $H(\lambda)$, $\partial_{\lambda}H(\lambda)$, and $A_{\mu}$’s are $k$-local operators with $k$ small and independent of the system size. More precisely, for a fixed $\lambda$, we assume that we can write $H(\lambda)=h_{1}+h_{2}+\cdots$, $\partial_{\lambda}H(\lambda)=l_{1}+l_{2}+\cdots$, and $A_{\mu}=a_{\mu 1}+a_{\mu 2}+\cdots$, where $h_{i}$, $l_{i}$, and $a_{\mu i}$ are operators that act nontrivially on at most $k$ sites. Based on this assumption, we can expand the action $\mathcal{S}_{\lambda}^{(1)}[V(\bm{\alpha})]$ in Eq. (6b) as
$\mathcal{S}_{\lambda}^{(1)}[V(\bm{\alpha})]$ $=-\hbar^{-2}\sum_{\mu ij}\sum_{\nu i^{\prime}j^{\prime}}\mathop{\mathrm{tr}}\left\{[h_{i},a_{\mu j}][h_{i^{\prime}},a_{\nu j^{\prime}}]\right\}\alpha_{\mu}\alpha_{\nu}$
In this expansion, $h_{i},a_{\mu j}]$ is zero if the domain of $h_{i}$ does not overlap the domain of $a_{\mu j}$. Otherwise, $h_{i},a_{\mu j}]$ is at most an (2k1)(2k-1)-local operator. Furthermore, if the domains of $h_{i},a_{\mu j}]$ and $l_{i^{\prime}}$ do not overlap, the trace $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]l_{i^{\prime}}\}$ is proportional to the product $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]\}\mathop{\mathrm{tr}}(l_{i^{\prime}})$, which is zero due to $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]\}=0$. Otherwise, the operator $h_{i},a_{\mu j}]l_{i^{\prime}}$ is at most (3k2)(3k-2)-local, and thus the trace $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}]l_{i^{\prime}}\}$ is determined by at most (3k2)(3k-2)-local information. Similarly, the trace $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}][h_{i^{\prime}},a_{\nu j^{\prime}}]\}$ is either zero or dependent on at most (4k3)(4k-3)-local information. Thus, the action $\mathcal{S}_{\lambda}^{(1)}[V(\bm{\alpha})]$ consists of local terms except for the “const.” term, which does not affect the minimization.
This locality of $\mathcal{S}_{\lambda}^{(1)}[V(\bm{\alpha})]$ does not necessarily imply that the resulting coefficient $\alpha_{\mu}$ depends only on local information. Nevertheless, this is true if $\mathop{\mathrm{tr}}\{[h_{i},a_{\mu j}][h_{i^{\prime}},a_{\nu j^{\prime}}]\}=0$ for any $\mu\neq\nu$, in which case the determination of $\alpha_{\mu}$ is decoupled from the determination of $\alpha_{\nu}$. This situation arises, for example, for locally interacting spin systems with the one-body driving [27, 28].
Derivation of Eqs. (11), (12), and (14)
We derive Eqs. (11), (12), and (14), which are central to the general framework in Sec. II.3. We fix a single value of $\lambda$ and omit the $\lambda$ dependencies of quantities such as $H(\lambda)$, $\epsilon_{n}(\lambda)$, $|\phi_{n}(\lambda)\rangle$, $\Phi(\lambda)$, $\mathcal{S}^{(\mathcal{P})}_{\lambda}[V]$, and $\mathcal{P}_{\lambda}(x)$ for notational simplicity.
To derive the algebraic characterization of the exact AGP in Eq. (11), we first calculate $\partial_{\lambda}\bigl(|\phi_{n}\rangle\langle\phi_{n}|\bigr)$ and $\partial^{\prime}_{\lambda}\mathcal{P}(H)$. The derivative of the projector $|\phi_{n}\rangle\langle\phi_{n}|$ is
$\partial_{\lambda}\bigl(|\phi_{n}\rangle\langle\phi_{n}|\bigr)$
$=|\phi_{n}\rangle\langle\partial_{\lambda}\phi_{n}|+|\partial_{\lambda}\phi_{n}\rangle\langle\phi_{n}|$
$=\sum_{m}|\phi_{n}\rangle\langle\partial_{\lambda}\phi_{n}|\phi_{m}\rangle\langle\phi_{m}|+\sum_{m}|\phi_{m}\rangle\langle\phi_{m}|\partial_{\lambda}\phi_{n}\rangle\langle\phi_{n}|$
$=i\hbar^{-1}\sum_{m}\bigl(\Phi_{nm}|\phi_{n}\rangle\langle\phi_{m}|-\Phi_{mn}|\phi_{m}\rangle\langle\phi_{n}|\bigr),$ (36)
where the last equality uses the definition of $\Phi$ in Eq. (4) and $\langle\partial_{\lambda}\phi_{m}|\phi_{n}\rangle+\langle\phi_{m}|\partial_{\lambda}\phi_{n}\rangle=\partial_{\lambda}[\langle\phi_{m}|\phi_{n}\rangle]=0$. Using Eq. (36), we can calculate the $\lambda$-derivative of the fictitious Hamiltonian $\mathcal{P}(H)=\sum_{n}\mathcal{P}(\epsilon_{n})|\phi_{n}\rangle\langle\phi_{n}|$ as
@@ Appendix A Appendix to the general framework
$\quad+i\hbar^{-1}\sum_{nm}\mathcal{P}(\epsilon_{n})\bigl(\Phi_{nm}|\phi_{n}\rangle\langle\phi_{m}|-\Phi_{mn}|\phi_{m}\rangle\langle\phi_{n}|\bigr)$
@@
$\quad+i\hbar^{-1}\sum_{nm}\bigl[\mathcal{P}(\epsilon_{n})-\mathcal{P}(\epsilon_{m})\bigr]\Phi_{nm}|\phi_{n}\rangle\langle\phi_{m}|$
@@
Next, we show that the algebraic characterization in Eq. (11) is equivalent to the variational expression in Eq. (12), which extends the derivation of the conventional variational method in Refs. [27, 8]. The algebraic characterization says that $\partial^{\prime}_{\lambda}\mathcal{P}(H)-i\hbar^{-1}[\mathcal{P}(H),\Phi]$ is diagonal with respect to the Hamiltonian basis. The diagonal elements come solely from $\partial^{\prime}_{\lambda}\mathcal{P}(H)$ since $\mathcal{P}(H),\Phi]$ has no diagonal elements. Let $\Psi\coloneqq\sum_{n}[\partial^{\prime}_{\lambda}\mathcal{P}(H)]_{nn}|\phi_{n}\rangle\langle\phi_{n}|$ denote the diagonal part of $\partial^{\prime}_{\lambda}\mathcal{P}(H)$. Then, the algebraic characterization is equivalent to
@@
Finally, we derive the expansion of $\mathcal{S}^{(\mathcal{P})}[V]$ in Eq. (14). Using the expression of $\partial^{\prime}_{\lambda}\mathcal{P}(H)$ in Eq. (37), we can rewrite the operator inside the Hilbert–Schmidt norm as
@@
Since the Hilbert–Schmidt norm equals the sum of the squared absolute values of the matrix elements, the action is given by
@@
$\quad+\hbar^{-2}\sum_{\smash{nm}}\bigl[\mathcal{P}_{\lambda}(\epsilon_{m})-\mathcal{P}_{\lambda}(\epsilon_{n})\bigr]^{2}|\Phi_{mn}-V_{mn}|^{2},$ (42)
which proves Eq. (14).
Inclusion of a conserved charge
When the Hamiltonian has a conserved charge, our formulation can incorporate it to derive a wider variety of weighted actions. Let $J(\lambda)$ be an operator that satisfies $H(\lambda),J(\lambda)]=0$, and we expand it in the Hamiltonian basis as
Introducing an arbitrary bivariate degree-$K$ polynomial $\mathcal{P}_{\lambda}(x,y)=\sum_{k=0}^{K}\sum_{l=0}^{k}p_{l,k-l}(\lambda)x^{l}y^{k-l}$, we can use $\mathcal{P}_{\lambda}(H(\lambda),J(\lambda))$ as a fictitious Hamiltonian in our framework. We can then determine driving coefficients by minimizing a weighted action of the form,
$\qquad\qquad\qquad\times\bigl|[V-\Phi(\lambda)]_{mn}\bigr|^{2}\Bigr\}+\mathrm{const.}$ (44)
For example, let us consider that $J(\lambda)\equiv J$ is a parity operator independent of $\lambda$ with the eigenvalues $j_{n}(\lambda)\equiv j_{n}=\pm 1$. The eigenstates are then split into even-parity ($j_{n}=+1$) and odd-parity ($j_{n}=-1$) states. Let us also assume that the driving operator $A_{\mu}$ has even parity for all $\mu$, meaning that $JA_{\mu}J=A_{\mu}$. Then, the matrix element $V(\bm{\alpha})-\Phi(\lambda)]_{mn}$ is nonzero only when the states $m$ and $n$ have the same parity. In this case, we can use $\mathcal{P}_{\lambda}(H(\lambda),J)=[H(\lambda)-E_{\lambda}^{(K)}]^{K}(J+1)$ to prioritize the suppression of nonadiabatic transitions between even-parity states. With this choice, $\mathcal{P}_{\lambda}(\epsilon_{n}(\lambda),j_{n})$ is zero for any odd-parity $n$, and thus the weighted action only penalizes the nonadiabatic transitions between even-parity states.
Note that the conventional variational method on a conserved subspace of $H(\lambda)$ has already been considered for a specific system. Reference [30] studies the $p$-spin Hamiltonian on spin-1/2 systems, which conserves the total spin. The study compares the conventional variational method applied to the subspace of the maximum total spin with that applied to the whole Hilbert space, finding that the former outperforms the latter for a severely underparameterized ansatz. Since including $J(\lambda)$ in our proposal is considered a much generalized version of this specific result, we expect that the inclusion of $J(\lambda)$ would indeed improve the driving fidelity.
Efficient algorithm for the entire range of $\lambda$
We discuss an efficient algorithm to compute the driving protocol over the whole range of $\lambda$, assuming that the Hamiltonian has the form of Eq. (19). We first introduce some notation. Let $\tilde{f}_{1}(\lambda),\tilde{f}_{2}(\lambda),\dots$ be the list of all the possible $k$-fold products of the coefficients $\{f_{1}(\lambda),\dots,f_{\Gamma}(\lambda)\}$ with $k=0,\dots,K$. For example, when $K=2$ and $\Gamma=2$, the list is $(\tilde{f}_{1}(\lambda),\dots,\tilde{f}_{6}(\lambda))=(f_{1}(\lambda)^{2},f_{1}(\lambda)f_{2}(\lambda),f_{2}(\lambda)^{2},f_{1}(\lambda),f_{2}(\lambda),1)$. The order of the list does not matter. We define $d_{g}$ as the total number of $f_{\gamma}$’s contained in $\tilde{f}_{g}(\lambda)$. In the above example, the numbers are $(d_{1},\dots,d_{6})=(2,2,2,1,1,0)$. We then define $\lambda$-independent operators $\tilde{F}_{1},\tilde{F}_{2},\dots$ by the relation,
$H(\lambda)^{k}=\left(\sum_{\gamma=1}^{\Gamma}f_{\gamma}(\lambda)F_{\gamma}\right)^{k}\eqqcolon\sum_{g:\,d_{g}=k}\tilde{f}_{g}(\lambda)\tilde{F}_{g}$ (45)
for $k=0,\dots,K$. In other words, $\tilde{F}_{g}$ is defined as the terms in $H(\lambda)^{d_{g}}$ whose $\lambda$-dependence is given by $\tilde{f}_{g}(\lambda)$. In the example above, the operators are $(\tilde{F}_{1},\dots,\tilde{F}_{6})=((F_{1})^{2},F_{1}F_{2}+F_{2}F_{1},(F_{2})^{2},F_{1},F_{2},I)$.
Using these definitions, we can conveniently separate $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ into $\lambda$-dependent coefficients and traces of $\lambda$-independent operators. By combining Eqs. (10) and (45), the fictitious Hamiltonian is rewritten as $\mathcal{P}_{\lambda}(H(\lambda))=\sum_{g}p_{d_{g}}(\lambda)\tilde{f}_{g}(\lambda)\tilde{F}_{g}$. Inserting this expression into the definitions of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$ in Eq. (17), we get
$\!\!Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ $=-\sum_{gg^{\prime}}p_{d_{g}}(\lambda)p_{d_{g^{\prime}}}(\lambda)\tilde{f}_{g}(\lambda)\tilde{f}_{g^{\prime}}(\lambda)\tilde{Q}_{\mu\nu,gg^{\prime}},$ (46a)
$\!\!r_{\mu}^{(\mathcal{P})}(\lambda)$ $=i\hbar\sum_{gg^{\prime}}p_{d_{g}}(\lambda)p_{d_{g^{\prime}}}(\lambda)[\partial_{\lambda}\tilde{f}_{g}(\lambda)]\tilde{f}_{g^{\prime}}(\lambda)\tilde{r}_{\mu,gg^{\prime}},$ (46b)
where we define
$\tilde{r}_{\mu,gg^{\prime}}$ $\coloneqq\mathop{\mathrm{tr}}\left\{\tilde{F}_{g}[\tilde{F}_{g^{\prime}},A_{\mu}]\right\}.$ (47b)
These expressions allow efficient computation of $Q_{\mu\nu}^{(\mathcal{P})}(\lambda)$ and $r_{\mu}^{(\mathcal{P})}(\lambda)$. To calculate them over the entire range of $\lambda$, we only need to calculate the traces $\tilde{Q}_{\mu\nu,gg^{\prime}}$ and $\tilde{r}_{\mu,gg^{\prime}}$ once.
Appendix B Appendix to the ground-state evolution
Derivation of the properties of the weight
@@ Appendix B Appendix to the ground-state evolution
We derive the properties of the weight $w_{mn}^{(K)}$ for ground-state evolution in Sec. III.1. We omit the $\lambda$ dependencies of $\epsilon_{n}(\lambda)$, $w_{mn}^{(K)}(\lambda)$, $w_{n}^{(K)}(\lambda)$, and $E_{\lambda}^{(K)}$ for conciseness.
First, we derive the explicit expression of $w_{mn}^{(K)}$ in Eq. (22) as
@@
$=\sum_{s=0}^{K-1}\sum_{u=0}^{K-1}(\epsilon_{n}-E^{(K)})^{(K-1-s)+u}(\epsilon_{m}-E^{(K)})^{(K-1-u)+s}$
$=\!\sum_{\smash{s=-(K-1)}}^{K-1}\!(K-|s|)(\epsilon_{n}-E^{(K)})^{K-1-s}(\epsilon_{m}-E^{(K)})^{K-1+s},$ (48)
where we combined the terms with the same value of sus-u in the last equality.
@@
Inserting this lower bound into Eq. (48) and combining with Eq. (50), we obtain